OpenStudy (anonymous):

Hi all! Q:"Where does the normal line to x^2 -xy+y^2=3 at P(-1,1) intersect the ellipse a second time?"

OpenStudy (anonymous):

I found the normal line to be y=x+2 but I'm not sure what to do next

OpenStudy (phi):

that equation for the normal line does not sound right. Did you use implicit differentiation to find dy/dx ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so I broke it up

OpenStudy (anonymous):

chain rule for x^2 and y^2

OpenStudy (anonymous):

Turn into 2x and 2yy'

OpenStudy (anonymous):

product rule for xy

OpenStudy (phi):

d/dy (x^2 -xy+y^2=3 ) 2x - d/dy(xy) + 2y dy/dx = 0 d/dy (x y)= x dy/dx + y Did you get this 2x - x y' - y + 2y y' = 0

OpenStudy (anonymous):

sec

OpenStudy (anonymous):

I got a different middle result I'll type it

OpenStudy (anonymous):

so product rule of xy is d/dx x(y)-d/dx y(x)

OpenStudy (anonymous):

which is then 1*y-y'x

OpenStudy (anonymous):

but I have to apply the minus sign

OpenStudy (phi):

except you do not subtract. d(xy)= x dy + y dx

OpenStudy (anonymous):

ah so for product rule you add?

OpenStudy (anonymous):

ok gotcha

OpenStudy (phi):

see https://en.wikipedia.org/wiki/Product_rule

OpenStudy (anonymous):

so then the derivative at the point is 1 not -1

OpenStudy (phi):

you should get 2x - x y' - y + 2y y' = 0 solve for y' what do you get ?

OpenStudy (anonymous):

(-y-2x)/(2y+x)

OpenStudy (anonymous):

=y'

OpenStudy (phi):

?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

nvm need to rewrite this section, sec pls :)

OpenStudy (anonymous):

so, y'= (-y-2x)/(2y-x)

OpenStudy (anonymous):

is this correct?

OpenStudy (phi):

closer. \[ 2x - x y' - y + 2y y' = 0 \\ 2y y' - x y' = y-2x\\ (2y -x)y'= (y-2x)\\ y'= \frac{y-2x}{2y -x} \]

OpenStudy (phi):

now evaluate at P(-1,1)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3/3 =1

OpenStudy (anonymous):

so the slope of the normal line is -1/1 =-1 then?

OpenStudy (phi):

yes. now use point- slope formula y - y0= m( x - x0) where (x0,y0) is (-1,1)

OpenStudy (anonymous):

y-1=-(x+1)

OpenStudy (anonymous):

or y=-x

OpenStudy (phi):

simplify to y = -x now you want to find where this line intersects the ellipse. sub in (-x) for y into x^2 -xy+y^2=3 and solve for the two x values

OpenStudy (anonymous):

aha I see

OpenStudy (phi):

one of the x values will be -1 (because -1,1 intersects the curve)

OpenStudy (anonymous):

then it should be 1 right?

OpenStudy (phi):

once you find the other x value, y = -x (or use the ellipse curve, to find y.. but that is harder)

OpenStudy (anonymous):

ok

OpenStudy (phi):

with y= -x in x^2 -xy+y^2=3 you get x^2 + x^2 + x^2 = 3 3x^2 = 3 x^2 = 1 x= ±1 so x=1 is the other x value.

OpenStudy (anonymous):

yep just finished, so the point is 1,-1

OpenStudy (phi):

Here is a graph of your problem

OpenStudy (anonymous):

ah ok, cool, nice to visualize. Thanks a lot for your help! :)